OpenStudy (anonymous):

if AB=2x-5 and BC=x+1 then find BC

OpenStudy (diyadiya):

Is there a figure?

OpenStudy (anonymous):

one moment drawing

OpenStudy (diyadiya):

Ok :)

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry

OpenStudy (diyadiya):

Can you label it? A ,B & C ?

OpenStudy (anonymous):

a left corner b at teop and c right corner

OpenStudy (anonymous):

d in the middle

OpenStudy (anonymous):

finding like values I think its 7 but not sure

OpenStudy (anonymous):

2x-5=x+x

OpenStudy (anonymous):

not right did not come thru right

OpenStudy (anonymous):

2x-5=x+1

OpenStudy (diyadiya):

So you have to find x ?